OpenStudy (anonymous):

7|x+1|+8=−3|x+1|+2 Help Please

hero (hero):

7|x+1|+8=−3|x+1|+2 7|x + 1| + 3|x + 1| = 2 - 8 (7 + 3)|x + 1| = -6 10|x + 1| = -6 |x + 1| = -6/10 The absolute value of a number can never be negative.

OpenStudy (anonymous):

Thank Youu(: